RHIC Run-10 Dry Run Plan
As of:
    October 6, 2009   



RHIC Dry Run (November 16 - 20, 2009)

Schedule:
Monday, November 16
Infrastructure
meeting 16:00 in MCR
Tuesday, November 17
Operations & Systems
meeting 16:00 in MCR
Wednesday, November 18
Systems
meeting 16:00 in MCR
Thursday, November 19
Instrumentation
meeting 16:00 in MCR
Friday, November 20
Contingency/Repeats
meeting 16:00 in MCR


Please check if:
- your name is assigned to the appropriate system
- your system is listed
- your system has or does not have changes
- all appropriate names are associated with your system
- if you are available that day or need to switch to another day
- if you have enough time


Pre-Dry Run Plans
October 1-9, FEC's, Servers, MCR Consoles, Permit link, event link's, remote scopes and other systems turn on.
During October all systems listed below will be brought on and tested individually, by the responsible contacts, as given below. Final check and testing will be performed during the dry run. By that time all systems will have been tested at least once, and any problems will have been fixed.


Day 1 (Infrastructure)


System Contact(s) comments status
changes
Elog
Todd, Ila

OK Nov 16
N
Event Link (& Display)
Todd, controls
 
OK Nov 16
N
Timing
Rob H., Rob M., Todd S.
beam synch, RHIC events, ...
OK Nov 16

BLAM
Jonathan L., Greg M., Leif A.
test
OK Nov 16-17
?
Alarms
Greg M., Seth N.
 
OK Nov 16
N
tape
Nick K., Greg M.

mostly OK Nov 16
Y
Database
Wenge F.
 
mostly OK Nov 16
N
RHIC Configuration Control
Greg M., Angelika
check status
OK Nov 16
Y
RF Configuration


In flux Nov 16 Y
Operations GPMs, StartUp
operations, Angelika, Greg M.
ramp, magnets, beam current, fill pattern, SpecMan, etc.
OK Nov 16
?
Logger
Roger L.
test
OK Nov 16
N
LogView
Ted
test
OK Nov 16
Y
Abort kickers
Leif A.
check out
OK Nov 18
N
RHIC Injection (& kickers)
Wolfram F.
event generation and detection, software changes?, check out

N
Web Forms
 
Ring Entry, Action Please, RS LOTO, etc.
RS LOTO OK, Ring Entry OK
Action Please OK, Trouble Reports OK

N
QLI Diagnostics
 
 
In flux Nov 16
Y
Cryo Logging
 
 
OK Nov 16 (on web)
 
Vacuum Logging
 
Booster, BtA, AGS, AtR, RHIC
OK Nov 16 (on web)
 
Call in Lists
 
started and up to date
OK Nov 20
 
BERT
Greg M., Seth
test
OK Nov 16 (no changes)
N
Print Utility
 
Updated for all new logs. Works?
OK Nov 16 (with new elogs)
 


 Day 2 (Operations & Systems)


System
commissioner
comments
status
changes
Permit Link
Greg M.
check out
OK Nov 18
N
Post Mortem
Jonathan L.
check out
Not tested yet Nov 20
N
Green Synchro
Jim J., Vincent
 
OK Nov 17
Y
Ramp
Steve T.
Defined and available default to applications?
OK Nov 18
Y
Ramping systems
& RHIC online model
Al, Steve, Guillaume, Todd
WFG manager, online model ...
OrbitCalc, online-offline interface ...
OK Nov 18
Y
Ramp Editor
Steve T, Todd, Ted
test
OK Nov 18
Y
RHIC Chromaticity
Steve T
test
OK Nov 17
?
ProfileDisplay
mcr
working? BtA
OK Nov 19
N
AGS Orbit Display
Leif A.
working?
OK Nov 19
Y ?
AGS Orbit Control
Jonathan L.
 
OK Nov 19
Y ?
Loss Monitors &
Thresholds
Mei, Seth
RHIC, AtR, injection data
OK Nov 17
N
AGS Tune Meter
Leif A., Jonathan L.
 
OK Nov 17
Y ?
AGS Main Magnet Control
 
Working?
OK Nov 20
N
AGS Optics Control
Keith Z.
Working?
Unknown Nov 20
N
AGS Gamma Jump
Keith Z.
Working?
Unknown Nov 20
N
AGS Online Model
Vincent S.
Working?
OK Nov 20
Y
AtR
Vincent S., Nick T.
setpoints for Au, AtR managers, AtR orbit, BPMs, loss monitors, emittance measurements (automatic measurements, records?)
OK Nov 17
?
Drift Compensation
Al M., Larry H.
 
OK Nov 18
Y
BBQ
Kevin M., Michiko M., Michelle W.
test, check out
OK Nov 17
Y
tune feedback
Al M., Michiko M.
test, check out
OK Nov 19
Y
chromaticity feedback
Al M., Michiko M., Steve T.
test, check out
OK Nov 20
Y
BTF
Wolfram F.
test, check out
OK Nov 17
Y
BPMs
Rob H., Rob M., Todd S.
managers, timing
OK Nov 18
Y
AGS IPM
Steve J.
working?
OK Nov 18
N

                

   Day 3 (Systems)

System
Commissioner
comment
status
changes
RHIC OrbitDisplay
Vadim, Todd
test
OK Nov 18
?
Orbit Correction
Vadim
test (new features?), ramp correction ...
OK Nov 18
Y
Tune Meter (Artus)
Angelika D., Rob M., Jonathan L.
test new software
OK Nov 18
Y
Injection Damper
Angelika D., Rob M.
check out
OK Nov 18
N
Plotfill, FDAview
Todd
test
OK Nov 18
Y
collimators
Angelika, Wenge F., Guillaume
test, start from sequencer
OK Nov 18
Y
gap cleaner
Angelika, Wenge
test, start from sequencer
OK Nov 18
Y
LISA
Ted, Angelika
test new features
OK Nov 18
Y
Tune Display
Ted
test
OK Nov 18
Y
Jump Quads, Jump Watch
Christoph
test
OK Nov 16
N
10 Hz Feedback
Rob M., Rob H.
status?, check out
OK Nov 25, magnets in/PS not
Y
Skew Mod, Coupling Angle
Yun, Joanne
check out
OK Nov 19
?
DQmin
Joanne
test
OK Nov 18
?
RhicLumi
Seth, Angelika
test
Unknown Nov 20
Y
Store Automation
Greg M., Angelika
status?
Unknown Nov 20
Y?
IR Bump
Fulvia, Sev
test
Comm problems Nov 19
N
RF, applications
Kevin S., Tom H.
check out
Unknown Nov 20
Y
Longitudinal Quad. Feedback
Kevin S.
ready?
Unknown Nov 20
Y
Yellow Dipole Radial Feedback
Carl S.
ready?
Unknown Nov 20
Y
Stochastic Cooling
Mike B., Mike B.
tests?
Some testing OK Nov 19
Y

                

   Day 4 (Instrumentation)

System
Commissioner
comment
status
changes
WCM
Roger L
test, new scopes, ethernet
OK Nov 20
Y
IPM
Roger C, Steve T
test application
OK Nov 19
?
Bunch Int. Monitor
Roger L., Wolfram F.
 
Unknown Nov 20
Y ?
RHIC vacuum, pressure display
Dick H, Haixin, Seth
test (application)
Unknown Nov 20
N
e- detectors
SY, Angelika, ...?
check out, test
Unknown Nov 20
?
AC dipole
Mei, Peter O., G. Wang
work test
Unknown Nov 20
N
transition monitor (button pickup)
Mike Blasko, Angelika, Roger L
test application
Being upgraded Nov 25, in progress
Y
Schottky
Roger L., Michelle W.
ready
OK Nov 19
Y
AGS polarimeter
Haixin, Seth
test polarimeter hardware and software in case of pp
OK Nov 18
Y
Coherence Monitor
 
 
OK Nov 18
 
ORM
Todd  
OK Nov 20
N


Please contact Kevin (kbrown@bnl.gov or x4409) or Todd (satogata@bnl.gov or x5826) for more information or comments.